# Cluster Synchronization via Graph Laplacian Eigenvectors

### v1

@@ IV.2 Spectral Cluster Synchronization With Phase Lag
$\dot{\theta_{i}}=\omega_{i}-\sigma\sum_{j\in V}A_{ij}\sin(\theta_{i}-\theta_{j}+\beta_{ij})$ (8)
@@
$\dot{\alpha}_{r}=\omega^{(r)}-\sigma\sum_{a\in E(G)}W_{aa}e_{a}^{(r)}\sin\left(\sum_{s>0}e_{a}^{(s)}\alpha_{s}(t)+\beta_{a}\right)$ (9)
The linearized equilibrium analysis reveals that close to equilibrium, and with small $\beta_{a}$, we get the following limit for the coefficients.
$\alpha_{r}^{\infty}=\frac{\omega^{(r)}-\sigma\sum_{a\in E(G)}W_{aa}e_{a}^{(r)}\beta_{a}}{\sigma\lambda_{r}}$
Let’s consider a couple examples to put this equation in context.
Example 1: Suppose the Kuramoto Sakaguchi model was placed on a network with an AEP, where intra cluster edges exhibit phase lag $\beta_{a}>0$, and inter cluster edges have negligible phase lag with $\beta_{a}\approx 0$. As before, natural frequencies are constant within clusters. In this case it follows that the equilibrium for all coefficients is.
$\alpha_{r}^{\infty}$ $=\frac{\omega^{(r)}-\sigma\sum^{\text{intra}}_{a}W_{aa}e_{a}^{(r)}\beta_{a}}{\lambda_{r}\sigma}$
Then structural coefficients for our AEP approach the limit $\frac{\omega^{(r)}}{\lambda_{r}\sigma},$ whereas for the non structural coefficients approach $-\frac{\sum^{\text{intra}}_{a}W_{aa}e_{a}^{(r)}\beta_{a}}{\lambda_{r}}.$ So in this example we see the structural coefficients approach the same values we would expect in a system without phase lag, exhibiting an isolation of the information for clustering behavior. Then the non structural coefficients, which dictate our distance from full cluster synchronization, are excited and reach their own equilibrium through an expression of the intra cluster phase lag and edge weights.
@@
where we define $\beta^{(r)}=e^{(r)}\cdot\vec{\beta}$. This is the alignment of the $r$th down edge Laplacian eigenvector with the phase offset vector. Assuming this network has an AEP and that the natural frequencies are constant within clusters, recall that the down edge Laplacian eigenvectors can each be interpreted as providing an edge value assignment. Now, considering that $\vec{\beta}$ is a linear combination of the vectors $e^{(r)}$ corresponding to the structural eigenvectors. The previous condition is equivalent to the inter-cluster edges between any pair of clusters being equal and the intra-cluster edges vanishing. In this case, for structural coefficients, the equilibrium state looks as above, but by orthogonality, the nonstructural coefficients have $\beta^{(r)}=0$. Thus, for nonstructural eigenmodes, $r$, we get $\alpha^{\infty}_{r}=0$, which means we can recover cluster synchronization in case of phase lag. Intuitively, our example requires the phase lag to respect the vertex partition. Moreover, when we allow edge weights to vay but keep a constant phase lag across all edges, the equilibrium state for the coefficients becomes
@@ V Discussion
The first consideration is the proximity to synchronization, as transient behavior can disrupt the change of basis. The phases of our oscillators are in $\mathbb{R}$ mod $2\pi$, but the basis of eigenvectors represents vectors in $\mathbb{R}$. Thus, it is possible for transient behavior on the way to synchronization to result in phase offsets in multiples of $2\pi$ between synchronized oscillators (a non trivial winding cell). These are equivalent phases mod $2\pi$, but not in the eigenvector basis, which results in the activation of spurious modes being necessary to describe the differences. This non-equivalence in the eigenvector basis means that we lose interpretability when the initial dynamics allow for some oscillators to get a full rotation ahead or behind. If this occurs at some point before synchronization, one could re-zero the phase vector closer to the synchronization to regain the interpretability of the decomposition thereafter. However, this property suggests that instead of transient dynamics or chaotic system wide behavior, the eigenvector basis is most suited to studying synchronization patterns, or fully clustered dynamics, as we have in this paper.
Our work has a clear relationship to spectral clustering techniques and the study of cluster detectability in networks. It is easy to show, for example, that a stochastic block matrix (SBM) model does approach an AEP as the number of vertices in the graph approaches infinity. The convergence to an AEP follows from showing the null expectation of the values in $E$, and the law of large numbers (See the appendix C for more of the proof).
@@
In the positive limit of $n$, random graphs sampled from the SBM model approach AEPs (or more precisely, EPs).
@@
This relates our notion of a QEP to the previously developed model of stochastic AEPs, which describes simple graphs which are almost equitable in expectation [10].
@@
Recently, structural eigenvectors of the Bethe Hessian have been used for spectral clustering algorithms in sparse networks [27]. We note that this matrix is a deformed Laplacian, sharing the same eigenvectors, but rescaling the eigenvalues to be between $-1$ and $1$. Thus one can study the same structural eigenvectors in relation to our spectral clusters [19] using 1 dimensional clustering [28] along individual eigenvectors or $k-$means on a block of eigenvectors to separate community structure.
@@
We note that graphs with high symmetry can allow for duplicate AEPs, which manifest in the eigenbasis with repeated eigenvalues. These repeated eigenvalues allow for variability in the choices of eigenvector representatives. In this case eigenvector representatives of one associated AEP may not have uniform values within clusters. This slightly changes how the results are interpreted, but no matter the choice of eigenvector representative, the sub basis of structural eigenvectors is still formed, allowing for the dimensionally reduced description of clustered dynamics. This kind of situation may lead to difficulty with detectability using 1-dimensional clustering methods because the choices of eigenvector representatives affect which group structure is displayed. However, in practice this effect doesn’t have a substantial impact as most networks don’t exhibit such high regularity, and this is even rarer for weighted networks.
@@ VI Conclusion
We expand on how the eigenbasis of the graph Laplacian encodes information fundamental to the synchronization of network models, and can be used to reformulate the Kuramoto model’s equations. This theoretical advancement reveals how the structure encoded in these eigenvectors provides direct insight into cluster synchronization and clustered dynamics, particularly when examining network vertex partitions into synchronizable clusters. Our framework recontextualizes AEP cluster synchronization as the dynamics of structural eigenvectors of the graph laplacian. We introduce quasi-equitable partitions (QEPs) as an approximate generalization of AEPs, allowing for deviations from the AEP spectrum. We prove that the spectral properties we establish for AEPs extend to QEPs when variations are sufficiently small, and we characterize the stability of these results as QEPs diverge from AEPs.
Furthermore, we demonstrate the versatility of our spectral approach by extending it to variants of the Kuramoto model, including networks with weighted edges and asymmetric phase lag. By unifying multiple theoretical frameworks for studying synchronization, our work not only bridges existing approaches but also opens new avenues for investigating complex network dynamics. This unified spectral perspective provides a promising foundation for future research in networked dynamical systems and their synchronization properties.
@@ Acknowledgements.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2235204. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Coefficient Form of Kuramoto
We show the derivation of equations 9 (and 2 by extension) onto weighted networks with antisymmetric phase lag, extending the equations as defined in [21]. Our antisymmetric phase lag assumption means that the phase offset adheres to $\beta_{ij}=-\beta_{ji}$. As we saw in section II.1, the graph Laplacian of a weighted graph can be written as $L=BWB^{T}$, where $W$ is a diagonal matrix of positive edge weights, and $B$ is the signed incidence matrix. As we saw with the unweighted Laplacian, there is an edge specific counterpart with which it shares a spectrum. Consider
@@
The last step turns the equation into a vector equation, and the sum from one over neighbors to one over incident edges. We denote the $a$th column of $B$ with $B_{a}$. This column corresponds to an edge, and contains a single $1$ and $-1$ to denote the choice of edge orientation. Note that two of these vectors are present in the vector formulation. Since sine is an odd function, the negative value in the $B_{a}$ outside flips the orientation of the difference $v_{i}^{(s)}-v_{j}^{(s)}=B_{a}^{T}v^{(s)}$ to the one appropriate for that vertex. The antisymmetry of the phase offset comes into effect here. For an edge $(i,j)$ with the orientation for the signed incidence as written, and identified as $a$, we let $\beta_{a}=\beta_{ij}$. Then $\beta_{ji}=-\beta_{a}$, which allows this offset to agree with the sign flip inside the sine function. We may now isolate the coefficient using the orthogonality of the basis.
@@
Something to note is that the sine term in the original Kuramoto model is called the phase interaction function (PIF) [9], and all of our analysis would work with other PIFs as long as they are odd functions, with only mild adjustments if the first order taylor approximation differs.
Appendix B Single mode dynamics for weighted graphs

### v2

@@ IV.2 Spectral Cluster Synchronization With Phase Lag
$\dot{\theta_{i}}=\omega_{i}-\sigma\sum_{j\in V}A_{ij}\sin(\theta_{i}-\theta_{j}+\beta_{ij}).$ (8)
@@
$\dot{\alpha}_{r}=\omega^{(r)}-\sigma\sum_{a\in E(G)}W_{aa}e_{a}^{(r)}\sin\left(\sum_{s>0}e_{a}^{(s)}\alpha_{s}(t)+\beta_{a}\right).$ (9)
The linearized equilibrium analysis reveals that close to equilibrium, and with small $\beta_{a}$, we get the following limit for the coefficients:
$\alpha_{r}^{\infty}=\frac{\omega^{(r)}-\sigma\sum_{a\in E(G)}W_{aa}e_{a}^{(r)}\beta_{a}}{\sigma\lambda_{r}}.$
The interested reader can find a full derivation of these equations in appendix A. Let’s consider a couple examples to put this equation in context.
Example 1: Suppose the Kuramoto Sakaguchi model was placed on a network with an AEP, where intra cluster edges exhibit phase lag $\beta_{a}>0$, and inter cluster edges have negligible phase lag with $\beta_{a}\approx 0$. As before, natural frequencies are constant within clusters. In this case it follows that the equilibrium for all coefficients is:
$\alpha_{r}^{\infty}$ $=\frac{\omega^{(r)}-\sigma\sum^{\text{intra}}_{a}W_{aa}e_{a}^{(r)}\beta_{a}}{\lambda_{r}\sigma}.$
Then structural coefficients for our AEP approach the limit $\frac{\omega^{(r)}}{\lambda_{r}\sigma},$ whereas for the non structural coefficients approach $\frac{\sum^{\text{intra}}_{a}W_{aa}e_{a}^{(r)}\beta_{a}}{\lambda_{r}\sigma}.$ So in this example we see the structural coefficients approach the same values we would expect in a system without phase lag, exhibiting an isolation of the information for clustering behavior. Then the non structural coefficients, which dictate our distance from full cluster synchronization, are excited and reach their own equilibrium through an expression of the intra cluster phase lag and edge weights.
@@
where we define $\beta^{(r)}=e^{(r)}\cdot\vec{\beta}$. This is the alignment of the $r$th down edge Laplacian eigenvector with the phase offset vector. Assuming this network has an AEP and that the natural frequencies are constant within clusters, recall that the down edge Laplacian eigenvectors can each be interpreted as providing an edge value assignment. Now suppose that $\vec{\beta}$ is a linear combination of the vectors $e^{(r)}$ corresponding to the structural eigenvectors. This requires equality in the antisymmetric phase-lag between pairs of clusters, with respect to the direction, and zero phase-lag within clusters. In this case, for structural coefficients, the equilibrium state looks as above, but by orthogonality, the nonstructural coefficients have $\beta^{(r)}=0$. Thus, for nonstructural eigenmodes, $r$, we get $\alpha^{\infty}_{r}=0$, which means we can recover cluster synchronization in case of phase lag. Intuitively, our example requires the phase lag to respect the vertex partition. Moreover, when we allow edge weights to vary but keep a constant phase lag across all edges, the equilibrium state for the coefficients becomes
@@ V Discussion
The first consideration is the proximity to synchronization, as transient behavior can disrupt the change of basis. The phases of our oscillators are in $\mathbb{R}$ mod $2\pi$, but the basis of eigenvectors represents vectors valued in $\mathbb{R}$. Thus, it is possible for transient behavior on the way to synchronization to result in phase offsets in multiples of $2\pi$ between synchronized oscillators (a non-trivial winding cell). These are equivalent phases mod $2\pi$, but not in the eigenvector basis, which results in the activation of spurious modes being necessary to describe the differences. This non-equivalence in the eigenvector basis means that we lose interpretability when the initial dynamics allow for some oscillators to get a full rotation ahead or behind. If this occurs at some point before synchronization, one could re-zero the phase vector closer to the synchronization to regain the interpretability of the decomposition thereafter. However, this property suggests that instead of chaotic system wide behavior, the eigenvector basis is most suited to studying synchronization patterns, or fully clustered dynamics, as we have in this paper.
Our work has a clear relationship to spectral clustering techniques and the study of cluster detectability in networks. It is easy to show, for example, that a stochastic block model (SBM) does approach an AEP as the number of vertices in the graph approaches infinity. The convergence to an AEP follows from showing the null expectation of the values in $E$, and the law of large numbers (See the appendix D for more of the proof).
@@
In the positive limit of $n$, there exists a natural vertex partition for sampled graphs, given by the SBM block structure, for which the deviation from the AEP condition (or more precisely, EP) vanishes in the large-size limit.
@@
This relates our notion of a $\delta$-QEP to the previously developed model of stochastic AEPs, which describes simple graphs which are almost equitable in expectation.Schaub, Li, and Peel (2023)
@@
Recently, structural eigenvectors of the Bethe Hessian have been used for spectral clustering algorithms in sparse networks.Duan and Mai (2020) We note that this matrix is a deformed Laplacian, sharing the same eigenvectors, but with rescaled eigenvalues. Thus one can study the same structural eigenvectorsKrzakala et al. (2013) in relation to $\delta$-QEP community structure using 1 dimensional clusteringPatania, Allard, and Young (2023) along individual eigenvectors or $k$-means on a block of eigenvectors to separate community structure.
@@
We note that graphs with high symmetry can allow for duplicate AEPs, which manifest in the eigenbasis with repeated eigenvalues. These repeated eigenvalues allow for variability in the choices of eigenvector representatives. In this case eigenvector representatives of one associated AEP may not have uniform values within clusters. This slightly changes how the results are interpreted, but no matter the choice of eigenvector representative, the sub-basis of structural eigenvectors is still formed, allowing for the dimensionally reduced description of clustered dynamics. This kind of situation may lead to difficulty with detectability using 1-dimensional clustering methods because the choices of eigenvector representatives affect which group structure is displayed. However, in practice this effect does not have a substantial impact as most networks don’t exhibit such high regularity, and this is even rarer for weighted networks.
While the focus of this work is theoretical, the spectral framework developed here has direct implications for the analysis of real-world systems where synchronization plays a functional role. In neuroscience, for example, both structural and functional brain networks exhibit modular and near-symmetric organization, which can support cluster synchronization across regions. Studies have shown that Laplacian eigenmodes align with empirically observed patterns of resting-state brain activity,Atasoy, Donnelly, and Pearson (2016); Deco, Jirsa, and McIntosh (2011); Cabral et al. (2011) suggesting that AEP- or $\delta$-QEP-based reductions could provide mechanistic insight into these patterns.
An exciting direction for future work involves extending this spectral approach from vertex dynamics to edge dynamics. Recent work by Faskowitz et al. has highlighted the role of edge-centric representations in brain connectivity,Faskowitz et al. (2020) and Maria Pope and colleagues have explored dynamic synchronization in networks where edge properties (e.g., weights or phase lags) change over time in response to activity.Pope et al. (2023) Adapting our framework to such edge-based models may provide new tools for capturing coordination at finer spatiotemporal resolutions and for modeling dynamic reconfiguration in systems where edge dynamics play an active role in shaping global synchrony.
Similarly, in power grid networks, synchronization of coupled oscillatory units (e.g., generators and loads) is essential for operational stability. The grid’s underlying modular structure often gives rise to approximate symmetries that could be captured by $\delta$-QEPs.Chow (2013) Previous work has used spectral and symmetry-based approaches to study synchronization and stability in such systems,Dorfler and Bullo (2012); Motter et al. (2013); Nishikawa and Motter (2015) and our framework may provide an additional low-dimensional tool for modeling and control in this context. A natural next step is to integrate this framework into empirical studies of these systems, further bridging the gap between network structure and dynamic behavior.
In general, many real-world systems, such as biological or social networks, include inhibitory or antagonistic interactions, naturally leading to signed networks. The current framework assumes all edge weights are strictly positive; this assumption is primarily made to preserve the positive semi-definiteness of the graph Laplacian and to ensure convergence properties of the Kuramoto dynamics. Extending the spectral framework to accommodate such systems would require analyzing the Laplacian of signed graphs. This introduces new challenges, such as the possibility of negative eigenvalues and less predictable anti-synchronization behavior. We see this as an important avenue for future research, particularly in modeling competitive interactions and studying how inhibitory coupling affects the formation and stability of clustered dynamics.
@@ VI Conclusion
Although our dynamical model is defined using the adjacency matrix, we emphasize that the Laplacian eigenvectors provide a particularly suitable basis for analyzing synchronization dynamics. This is because the Laplacian as an operator captures diffusion over the network, and its eigenbasis naturally encodes both global coherence (through the zero eigenmode) and modular organization (through low-lying nontrivial eigenvectors). The Kuramoto model is a diffusive system and its linearization can be expressed directly in terms of the Laplacian. More generally, the Laplacian spectrum reveals how perturbations propagate and decay across the network, offering a principled way to identify and interpret emergent synchronizing clusters. Thus, using the Laplacian as the analytic backbone of our framework ensures that we capture both the structural and dynamical factors shaping clustered synchronization.
We expand on how the eigenbasis of the graph Laplacian encodes information fundamental to the synchronization of network models, and can be used to reformulate the Kuramoto model’s equations. This theoretical advancement reveals how the structure encoded in these eigenvectors provides direct insight into cluster synchronization and clustered dynamics, particularly when examining network vertex partitions into synchronizable clusters.
Our framework recontextualizes AEP cluster synchronization as the dynamics of structural eigenvectors of the graph laplacian. We introduce $\delta$ level quasi-equitable partitions ($\delta$-QEPs) as an approximate generalization of AEPs, allowing for deviations from the AEP spectrum. We prove that the spectral properties we establish for AEPs extend to $\delta$-QEPs when variations are sufficiently small, and we characterize the stability of these results as $\delta$-QEPs diverge from AEPs.
Furthermore, we demonstrate the versatility of our spectral approach by extending it to variants of the Kuramoto model, including networks with weighted edges and antisymmetric phase lag. By unifying multiple theoretical frameworks for studying synchronization, our work not only bridges existing approaches but also opens new avenues for investigating complex network dynamics. This unified spectral perspective provides a promising foundation for future research in networked dynamical systems and their synchronization properties.
@@ Acknowledgements.
We thank the Reviewers and the Editor for their many helpful suggestions and comments. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2235204. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
Author Declarations
The authors have no conflicts to disclose
Data Availability Statement
The data and code that support the findings of this paper, and recreate the figures, are openly available on github with the following link https://github.com/TobiasTimofeyev/Spectral_KuramotoTimofeyev (2025).
@@ Appendix A Coefficient Form of Kuramoto
We show the derivation of equations 9 (and 2 by extension) onto weighted networks with antisymmetric phase lag, extending the equations as derived by Kalloniatis.Kalloniatis (2010) Our antisymmetric phase lag assumption means that the phase offset adheres to $\beta_{ij}=-\beta_{ji}$. As we saw in section II.1, the graph Laplacian of a weighted graph can be written as $L=BWB^{T}$, where $W$ is a diagonal matrix of positive edge weights, and $B$ is the signed incidence matrix. As we saw with the unweighted Laplacian, there is an edge specific counterpart with which it shares a spectrum. Consider
@@
The last step turns the equation into a vector equation, and the sum from one over neighbors to one over incident edges. We denote the $a$th column of $B$ with $B_{a}$. This column corresponds to an edge, and contains a single $1$ and 1-1 to denote the choice of edge orientation. Note that two of these vectors are present in the vector formulation. Since sine is an odd function, the negative value in the $B_{a}$ outside flips the orientation of the difference $v_{i}^{(s)}-v_{j}^{(s)}=B_{a}^{T}v^{(s)}$ to the one appropriate for that vertex. The antisymmetry of the phase offset comes into effect here. For an edge $\{i,j\}$ with the orientation for the signed incidence as written, and identified as $a$, we let $\beta_{a}=\beta_{ij}$. Then $\beta_{ji}=-\beta_{a}$, which allows this offset to agree with the sign flip inside the sine function. We may now isolate the coefficient using the orthogonality of the basis.
@@
Something to note is that the sine term in the original Kuramoto model is called the phase interaction function (PIF),Breakspear, Heitmann, and Daffertshofer (2010) and all of our analysis would work with other PIFs as long as they are odd functions, with only mild adjustments if the first order taylor approximation differs.
Appendix B Transient Clustering
We begin with the proof of proposition 5.
Proof.
Let $G$ be a connected graph with an AEP, $\pi$. Suppose $\pi$ is sufficiently assortative in that the $p\geq 1$ smallest eigenvalues of the graph Laplacian, $L$, are associated to corresponding structural eigenvectors. Assume there are no repeated eigenvalues. Consider a Kuramoto model on $G$ with uniform natural frequencies, that is $\vec{\omega}=c\mathbb{1}_{n}$ for some constant $c$.
Recall the explicit solution for the linearized equations, 4. Since $G$ is connected it follows that the eigenvalue $\lambda_{0}=0$ of $L$ has multiplicity one, and $\vec{v}_{0}=\frac{1}{\sqrt{N}}\mathbb{1}_{n}$. By orthogonality it follows that $\omega^{(r)}=0$ for all $0<r\leq n$. Thus equation 4 reduces to
$\alpha_{r}(t)=\alpha_{r}(0)e^{-\sigma\lambda_{r}t}$
for all $0<r\leq n$. Since $\sigma$ is a constant, the positive eigenvalue distinguishes the relative decay of differing eigenmodes fromt their starting values. By assumption, for modes $0<r\leq p$, the eigenvectors are structural, and without repeated eigenvalues, are in the column space of $P^{\pi}$. Thus when these modes dominate the dynamics of $\vec{\theta}$, cluster synchronization is achieved. Notice that for $0<i\leq p$ and $p<j\leq n$,
$\frac{|\alpha_{i}(t)|}{|\alpha_{j}(t)|}=\frac{|\alpha_{i}(0)|}{|\alpha_{j}(0)|}e^{-\sigma(\lambda_{i}-\lambda_{j})t}.$
By assumption $\lambda_{i}<\lambda_{j}$, which implies that this ratio is growing with time for all such pairs of modes where $|\alpha_{i}(0)|,|\alpha_{j}(0)|>0$. Thus, naturally the eigenmodes with indices $0<i\leq p$ and non-zero initial condition will dominate with time.
Note that the choice of initial condition has a significant impact on the visibility of the clustered behavior, as the structural eigenmodes may only dominate once all coefficient magnitudes have become small. ∎
We note that with additional restrictions for repeated eigenvalues and eigenvectors, the above proof can straightforwardly be extended to the case of repeated eigenvalues. The clusters of the above emerge from a difference in the rates of decay. Since repeated eigenvalues create ambiguity in the choices of eigenvectors to span their induced subspace, a simple requirement that the choices of structural eigenvectors of the $p$ strictly smallest eigenvalues exist in the column space of $P^{\pi}$ is required.
Next we prove a lemma towards the proof of 6.
Lemma 12.
Consider two AEPs, $\pi_{1}$ and $\pi_{2}$, such that the former is a refinement of the latter. Then the column space of $P^{\pi_{2}}$ is contained in the column space of $P^{\pi_{1}}$.
Proof.
Let $w_{i}$ be the $i$th column of $P^{\pi_{2}}$. This has entries of $1$ for all vertex indices within cell $i$ of $\pi_{2}$. Since $\pi_{1}$ is a partition of the vertices, all the vertices of cell $i$ in $\pi_{2}$ are contained in some cell of $\pi_{1}$. Since $\pi_{1}$ is a refinement, $i$ is either a block in $\pi_{2}$ in its entirety, or subdivided into smaller blocks in $\pi_{1}$. In either case there exists some set of columns $u_{1},\dotsc,u_{p}$ such that $w_{i}=u_{1}+\dots+u_{p}$. Thus all the columns of $P^{\pi_{2}}$ are in the columns space of $P^{\pi_{1}}$ and we are done. ∎
Now, the proof of proposition 6 follows from the previous lemma and proposition 5.
Proof.
We assume that some low structural modes for each AEP in the hierarchy have nonzero initial condition. Under proposition 5, in time the eigenvalues with indices below $p_{k}$ will dominate the rest to whatever degree desired. By lemma 12, the combination of the $p_{i}$ lowest eigenvectors is in the column space of $P^{\pi_{i}}$, corresponding to the finest partition. Thus behavior dominated by the $p_{k}$ lowest eigenmodes corresponds to clustered behavior on $\pi_{k}$. However, proposition 5 may be applied again to the $p_{k-1}$ smallest modes, and the $p_{k-2}$ smallest modes, and so on. Thus, up to the choice of initial conditions, the system can experience cluster synchronization corresponding to each level of the AEP hierarchy, on the way to full system synchronization. ∎
Appendix C Single mode dynamics for weighted graphs